For any real numbers x and y ,coax =cosy ,implies =2nπ±y, where n belongs to Z 
In this solution sinx+y/2=0
How x+y/2=nπ?  please explain this steps

Grade:11

2 Answers

Aditya Gupta
5 years ago
you are asking how does sin(x+y/2)=0 imply x+y/2=nπ.
let  x+y/2=z
then sinz=0 but graph of sinx shows that it is zero if and only if z is an integral multiple of pi.
so z= npi
or  x+y/2=nπ
manantomar
5 years ago
 
you are asking how does sin(x+y/2)=0 imply x+y/2=nπ.
let  x+y/2=z
then sinz=0 but graph of sinx shows that it is zero if and only if z is an integral multiple of pi.
so z= npi
or  x+y/2=nπ
This is CORRECT